# latexText page 4 contains only floats

### v2

@@ 2.1.2 Metadata
Seismic intensity indicators include the Modified Mercalli Intensity (MI or MMI), Peak Ground Acceleration (PGA), Peak Ground Velocity (PGV) and Pseudo-Spectral Acceleration (PSA) at intervals of 0.3 s, 1.0 s, and 3.0 s. MI assesses earthquake impact by considering its effects on people, structures, and the landscape, with values ranging from 1 to 10, where higher numbers indicate more intense shaking and damage. PGA is a quantitative measure of the maximum ground acceleration during an earthquake, particularly relevant for short buildings (up to seven stories) USGS (Accessed: 2023-06-05a). PGV is an indicator of hazard for taller buildings USGS (Accessed: 2023-06-05a). PSA accounts for the hazard to buildings but is more closely aligned with how buildings respond to seismic forces at different frequencies, compared to peak ground motion parameters. These indices offer meaningful insights into earthquake severity and were obtained through the USGS website USGS (Accessed: 2023-06-05b). Additionally, the final condition of a structure can result not only from the primary earthquake but also from subsequent aftershocks. To account for the impact of aftershocks, we have included the top five Multipl in our metadata for all seismic intensity indicators.
SAR-derived parameters include geocoded single look complex (SLC) amplitude data and DPMs. Geocoded SLC amplitudes data provides information about back-scattered amplitude of the Earth’s surface in the microwave wavelength. In this study, we used VV (vertical transmit, vertical receive) and VH (vertical transmit, horizontal receive) polarization values derived from Sentinel-1 ground range detected (GRD) data. These polarizations reveal details about surface orientation and scattering properties. DPMs serve as indicators of surface changes, and are derived by detecting changes in coherence between pre- and post earthquake images. SAR data was accessed through the Alaska Satellite Facility (ASF) via NASA’s Earthdata Search API Earthdata (Accessed: 2023-06-05), while DPMs were extracted from the Jet Propulsion Laboratory’s data repository ARIA (Accessed: 2023-06-05).
@@ 2.2 Data Pre-Processing
An important step after collecting the imagery and the metadata, is to accurately map the ground truth of damage state of the building with the satellite imagery and it’s corresponding metadata. The ground truth damage state for the buildings was acquired from the Turkish Ministry of Environment, Urbanization and Climate Change Foursquare (Accessed: 2023-03-05). The damage severity in the dataset were classified into 4 damage classes, numbered 1-4, Slightly Damaged, Heavily Damaged, Needs to be demolished, and Collapsed. Mapping the ground truth with the dataset was done in two steps: mapping the ground truth to the building footprint, and then retrieving its corresponding meta data. In the first step, the nearest building footprint for each ground truth damage location (latitude, longitude) was identified by calculating spatial distances using an R-tree index Guttman (1984). Then, the data was filtered based on whether each ground truth damage location fell within its corresponding building footprint to ensure accurate mapping. In the second step, metadata was associated with the mapped building footprints by using the geometric centroid of each footprint as a query point. For contour shapefile data, such as seismic intensity indicators, interpolation was employed to estimate values at the query point. SAR-derived parameters and soil properties were extracted from raster GeoTIFF files. The raster data, originally in a global coordinate system measured in degrees (EPSG:4326), was reprojected into a local coordinate system measured in meters (EPSG:32637) to ensure precise alignment with the ground truth damage locations. This approach ensured that each building point was accurately associated with the relevant metadata.
@@
In addition to our primary dataset, we utilized another dataset developed by Yu et al. (2024) to train the model and ensure a fair comparison with other methods. We extracted the satellite imagery for this dataset following the methodology described above and mapped it to the corresponding metadata. This dataset includes five metadata features: damage proxy (DP) derived from Sentinel-1 and ALOS-2 PALSAR-2, ADI, NDBI, and PGA. The dataset has 5 damage classes numbered 0-4, No damage, Slightly Damaged, Heavily Damaged, Needs to be demolished, and Collapsed. The details of both datasets are summarized in the table 1.
@@
Name Turkey-L Turkey-S
@@
Region Turkey (All major cities) Kahramanmaraş City
@@ 2.3 QuakeMetaFormer
Levaraging the MetaFormer architecture proposed by Diao et al. (2022), illustrated in Figure 4, QuakeMetaFormer is a multi-input model that takes both image data and metadata as inputs optimized for post-earthquake PDA by accounting for extreme class imbalance. MetaFormer employs a hybrid approach where convolution is utilized to encode visual data, while transformer layers focus on combining this visual information with metadata, which is encoded using a non-linear embedding. The initial three stages of MetaFormer primarily utilize MBConv blocks, followed by the adoption of Relative Transformer blocks in the final two stages. To optimize computational efficiency, overlapping patch embedding is implemented for tokenizing the feature map and performing downsampling. The model configuration utilizes key hyperparameters listed in Table 2. The hyperparameters from the original MetaFormer paper were directly adopted from Diao et al. (2022).
@@
where $w_{t}$ represents the class weight for the true class, $p_{t}$ is the predicted probability of the true class, $\gamma$ is the focusing parameter, and $\text{CE}=-\log(p_{t})$ denotes the cross-entropy loss. We found that a gamma value of 4 worked best for our dataset, effectively focusing on hard-to-classify samples.
@@
Table 3: Various Metadata Embedding Results with Descriptions
Dataset Embedding F1 Score Description
Turkey-L MLP 0.49 Standard feed-forward stack
Turkey-L ResNorm 0.48 Residual connections with LayerNorm
Turkey-L DeepResNorm 0.48 Stacked ResNorm blocks
Turkey-L DeepMLP 0.47 Extended feed-forward network
Turkey-L Transformer 0.43 Standard Transformer encoder
Turkey-L LargeTransformer 0.43 2-layer Transformer encoder
Turkey-S MLP 0.56 Standard feed-forward stack
Turkey-S ResNorm 0.54 Residual connections with LayerNorm
Turkey-S DeepMLP 0.54 Extended feed-forward network
Turkey-S DeepResNorm 0.53 Stacked ResNorm blocks
Turkey-S Transformer 0.32 Standard Transformer encoder
Turkey-S LargeTransformer 0.23 2-layer Transformer encoder
Another modification we introduced focuses on enhancing the incorporation of metadata through different embedding strategies. We compared multiple approaches ResNorm, MLP, and Transformers. ResNorm combines residual connections with layer normalization, applying linear transformations and ReLU activations in a skip-connected fashion. In contrast, the MLP approach is a feed-forward stack of linear layers and activations followed by a final layer normalization. Meanwhile, the Transformer-based model leverages multi-head self-attention and feed-forward sublayers to encode metadata embeddings. We implemented two versions of each of these embeddings with deeper architectures. The 3 below presents the F1 scores for each dataset, along with brief explanations of the modifications made in the deeper variants.The results show that the MLP embedding strategy consistently achieves the highest F1 scores for both datasets, indicating that the feed-forward architecture performs robustly compared to the other methods. Additionally, while the deeper versions of ResNorm and MLP exhibit comparable performance to their standard counterparts, the Transformer-based approaches record noticeably lower F1 scores, particularly on the Quake-S dataset.
@@ 3 Experiments
In this section, we evaluate the performance of the proposed PDA framework by conducting four experiments: (i) impact of integrating metadata with satellite imagery on the overall model performance, (ii) comparing the performance of the proposed framework with two previously published state-of-the art machine learning frameworks, (iii) assessing the influence of individual metadata features on damage state prediction, and (iv) evaluating the effect of metadata in improving the trained model’s generalizability to different regions.
@@ 3.1 Impact of Adding Metadata on Performance: QuakeMetaFormer
Satellite imagery provides essential visual information for damage assessment, while earthquake-related metadata, such as PGA, SAR, VS30, etc., offers unique insights into event intensity and structural changes. However, relying on either data source individually often lacks the detail necessary for distinguishing between intermediate damage classes. The objective of this experiment is to evaluate how incorporating metadata with satellite imagery impacts the performance of the QuakeMetaFormer model.To achieve this, we conducted three experiments: (i) QIF-Turkey-L, where the QuakeImageFormer (QIF) was trained using only satellite imagery to establish baseline performance, (ii) QMF-Turkey-L, where the QuakeMetaFormer (QMF) model was trained using both satellite imagery and all the metadata, including aftershock seismic intensity indicators; and (iii) QMF-Turkey-L1, where the QuakeMetaFormer model was trained using both satellite imagery and metadata, but with the seismic intensity indicators for aftershocks removed, retaining only data from the main event of the largest magnitude. Similarly, two additional experiments, QIF-Turkey-S and QMF-Turkey-S, were conducted on the Turkey-S dataset. Comparing the results from these experiments allows us to understand the impact of adding metadata to satellite imagery in multiclass damage classification. Additionally, to evaluate the impact of metadata on the model’s ability to interpret imagery, we conducted four additional experiments. In these experiments, metadata was incorporated during training but masked during testing, simulating real-world scenarios where high-latency data, such as SAR, may not be readily available. Different models trained are summarized in Table 4. In each of our experiments, we utilize 85% of the data for training and reserve 15% for the testing set to evaluate model performance. Each experiment is discussed in detail in the following sections.
@@
QMF-Turkey-L QuakeMetaFormer (QMF) Turkey-L Yes No
QMF-Turkey-L-M QuakeMetaFormer (QMF) Turkey-L Yes Yes
QMF-Turkey-L1 QuakeMetaFormer (QMF) Turkey-L Yes No
QMF-Turkey-L1-M QuakeMetaFormer (QMF) Turkey-L Yes Yes
QIF-Turkey-L QuakeImageFormer (QIF) Turkey-L No N/A
QMF-Turkey-L1-M-DPM QuakeMetaFormer (QMF) Turkey-L Yes Yes (Only DPM)
QMF-Turkey-S QuakeMetaFormer (QMF) Turkey-S Yes No
QMF-Turkey-S-M QuakeMetaFormer (QMF) Turkey-S Yes Yes
QIF-Turkey-S QuakeImageFormer (QIF) Turkey-S No N/A
@@ 3.2 Comparison of QuakeMetaFormer and Machine Learning Frameworks
To evaluate the performance of our proposed framework against other state-of-the-art PDA frameworks for earthquake damage assessment, we compared different architectures trained on the same dataset. For this experiment, we trained the QuakeMetaFormer model (QMF-Turkey-S) on the Turkey-S dataset and compared its results with the best-performing machine learning model (random forest) as reported by Yu et al. (2024) on the same dataset. On scanning the literature, we found another paper, citeRao2023EarthquakeLearning that conducted a multiclass classification study for post-earthquake damage evaluation across regions impacted by 2015 Gorkha, 2017 Puebla, 2020 Puerto Rico, and 2020 Zagreb earthquakes. However, the datasets from Rao et al. (2023) were not accessible, so we report the average performance of our model across both our datasets and the average performance across all datasets in Rao et al. The experiments are summarized in Table 5.
@@
QMF-Turkey-S QuakeMetaFormer (QMF) Our Study
RF-Xiao Random Forest (RF) Yu et al. (2024)
@@ 3.3 Feature Importance Across Metadata
Various metadata provide different insights related to earthquake damage assessment. To assess the contribution of individual metadata features, we conducted a detailed analysis using SHapley Additive exPlanations (SHAP) Lundberg and Lee (2017). SHAP provides a robust method for understanding the output of machine learning models by perturbing each feature and measuring the resulting changes in prediction probabilities. These changes are then used to assign an importance score based on the feature’s contribution to the predictions. For this experiment, we applied SHAP to the QMF-Turkey-L1 model to evaluate the influence of various metadata inputs on damage state prediction. This analysis helps in understanding the impact of different earthquake-related indices on model performance. We also performed a class-wise analysis of metadata features to better understand the impact on different classes.

### v3

@@ 2.1.2 Metadata
Seismic intensity indicators include the Modified Mercalli Intensity (MI or MMI), Peak Ground Acceleration (PGA), Peak Ground Velocity (PGV) and Pseudo-Spectral Acceleration (PSA) at intervals of 0.3 s, 1.0 s, and 3.0 s. MI assesses earthquake impact by considering its effects on people, structures, and the landscape, with values ranging from 1 to 10, where higher numbers indicate more intense shaking and damage. PGA is a quantitative measure of the maximum ground acceleration during an earthquake, particularly relevant for short buildings (up to seven stories) USGS (Accessed: 2023-06-05a). PGV is an indicator of hazard for taller buildings USGS (Accessed: 2023-06-05a). PSA accounts for the hazard to buildings but is more closely aligned with how buildings respond to seismic forces at different frequencies, compared to peak ground motion parameters. These indices offer meaningful insights into earthquake severity and were obtained through the USGS website USGS (Accessed: 2023-06-05b). Additionally, the final condition of a structure can result not only from the primary earthquake but also from subsequent aftershocks. To account for the impact of aftershocks, we have included the top five earthquakes by intensity in our metadata for all seismic intensity indicators.
SAR-derived parameters include geocoded single look complex (SLC) amplitude data and DPMs. Geocoded SLC amplitudes data provides information about back-scattered amplitude of the Earth’s surface in the microwave wavelength. In this study, we used VV (vertical transmit, vertical receive) and VH (vertical transmit, horizontal receive) polarization values derived from Sentinel-1 ground range detected (GRD) data. These polarizations reveal details about surface orientation and scattering properties. DPMs serve as indicators of surface changes, and are derived by detecting changes in coherence between pre- and post earthquake images. These surface variations are utilized in buildings to identify structural changes, which in turn aids in damage assessment. SAR data was accessed through the Alaska Satellite Facility (ASF) via NASA’s Earthdata Search API Earthdata (Accessed: 2023-06-05), while DPMs were extracted from the Jet Propulsion Laboratory’s data repository ARIA (Accessed: 2023-06-05).
@@
Delays in data acquisition critically affects timely post‐disaster damage assessment, with acquisition timelines typically scaling with the size of area to be surveyed. Figure 4(a) and 4(b) shows progress of data acquisition over time, and spatial distribution across the study area respectively. Around 71% of the optical satellite data for the affected region were acquired within 1-day delay, reaching 100% by day 5. Also, Figure 4(b) suggests that data collection was prioritized for major cities near the epicenter, while remote areas and non-cities seems to experienced 3–5-day delays. The utility of the collected optical imagery is dependent on the weather conditions prevalent at the time as cloud cover may occlude buildings necessitating reacquisition. For metadata, SAR‐derived parameters (acquired Feb 11, 2023) incurred a 5-day delay, whereas USGS seismic intensity data are generally real‐time within the USA, but can be delayed 20–45 min for large earthquakes outside the US USGS Website Delay .
Figure 4: Optical Satellite Imagery Delays: (a) Acquisition over Time and (b) Spatial Distribution.
@@ 2.2 Data Pre-Processing
An important step after collecting the imagery and the metadata, is to accurately map the ground truth of damage state of the building with the satellite imagery and it’s corresponding metadata. The ground truth damage state for the buildings was acquired from the Turkish Ministry of Environment, Urbanization and Climate Change, Çevre, Şehircilik ve İklim Değişikliği Bakanlığı (CSB) in Turkish Foursquare (Accessed: 2023-03-05). These data were processed by Gece Yazılım and Yer Çizenler respectively. The damage severity in the dataset were classified into 4 damage classes, numbered 1-4, Slightly Damaged, Heavily Damaged, Needs to be demolished, and Collapsed. Mapping the ground truth with the dataset was done in two steps: mapping the ground truth to the building footprint, and then retrieving its corresponding meta data. In the first step, the nearest building footprint for each ground truth damage location (latitude, longitude) was identified by calculating spatial distances using an R-tree index Guttman (1984). Then, the data was filtered based on whether each ground truth damage location fell within its corresponding building footprint to ensure accurate mapping. In the second step, metadata was associated with the mapped building footprints by using the geometric centroid of each footprint as a query point. For contour shapefile data, such as seismic intensity indicators, interpolation was employed to estimate values at the query point. SAR-derived parameters and soil properties were extracted from raster GeoTIFF files. The raster data, originally in a global coordinate system measured in degrees (EPSG:4326), was reprojected into a local coordinate system measured in meters (EPSG:32637) to ensure precise alignment with the ground truth damage locations. This approach ensured that each building point was accurately associated with the relevant metadata.
@@
In addition to our primary dataset, we used another dataset developed by Yu et al. (2024), including the ground truth labels and metadata. We then extracted the corresponding satellite imagery using our methodology and mapped it to the metadata for model training. This dataset includes five metadata features: damage proxy (DP) derived from Sentinel-1 and ALOS-2 PALSAR-2, ADI, NDBI, and PGA. The dataset has 5 damage classes numbered 0-4, No damage, Slightly Damaged, Heavily Damaged, Needs to be demolished, and Collapsed. The details of both datasets are summarized in the table 1.
@@
Name Türkiye-L Türkiye-S
@@
Region Türkiye (All major cities) Kahramanmaraş City
@@ 2.3 QuakeMetaFormer
Leveraging the MetaFormer architecture proposed by Diao et al. (2022), illustrated in Figure 5, QuakeMetaFormer is a multi-input model that takes both image data and metadata as inputs optimized for post-earthquake PDA by accounting for extreme class imbalance. MetaFormer employs a hybrid approach where convolution is utilized to encode visual data, while transformer layers focus on combining this visual information with metadata, which is encoded using a non-linear embedding. The initial three stages of MetaFormer primarily utilize MBConv blocks, followed by the adoption of Relative Transformer blocks in the final two stages. To optimize computational efficiency, overlapping patch embedding is implemented for tokenizing the feature map and performing downsampling. The model configuration utilizes key hyperparameters listed in Table 2. The hyperparameters from the original MetaFormer paper were directly adopted from Diao et al. (2022).
@@
where $w_{t}$ represents the class weight for the true class, $p_{t}$ is the predicted probability of the true class, $\gamma$ is the focusing parameter, and $\text{CE}=-\log(p_{t})$ denotes the cross-entropy loss. We found that a gamma value of 4 worked best for our dataset, effectively focusing on hard-to-classify samples (see Table 3).
@@
Table 3: Comparison of Various Metadata Embeddings, Loss Functions, and F1 Scores
Dataset Embedding Loss Function F1 Score Description
Türkiye-L MLP MFL 0.49 Standard feed-forward stack
Türkiye-L MLP CEL 0.31 Standard feed-forward stack
Türkiye-L ResNorm MFL 0.48 Residual connections with LayerNorm
Türkiye-L DeepResNorm MFL 0.48 Stacked ResNorm blocks
Türkiye-L DeepMLP MFL 0.47 Extended feed-forward network
Türkiye-L Transformer MFL 0.43 Standard Transformer encoder
Türkiye-L LargeTransformer MFL 0.43 2-layer Transformer encoder
Türkiye-S MLP MFL 0.56 Standard feed-forward stack
Türkiye-S MLP CEL 0.34 Standard feed-forward stack
Türkiye-S ResNorm MFL 0.54 Residual connections with LayerNorm
Türkiye-S DeepMLP MFL 0.54 Extended feed-forward network
Türkiye-S DeepResNorm MFL 0.53 Stacked ResNorm blocks
Türkiye-S Transformer MFL 0.32 Standard Transformer encoder
Türkiye-S LargeTransformer MFL 0.23 2-layer Transformer encoder
CEL: Cross-Entropy Loss, MFL: Modified Focal Loss
Another modification we introduced focuses on enhancing the incorporation of metadata through different embedding strategies. We compared multiple approaches: ResNorm, MLP, and Transformers. ResNorm combines residual connections with layer normalization, applying linear transformations and ReLU activations in a skip-connected fashion. In contrast, the MLP approach is a feed-forward stack of linear layers and activations followed by a final layer normalization. Meanwhile, the Transformer-based model leverages multi-head self-attention and feed-forward sublayers to encode metadata embeddings. We implemented two versions of each of these embeddings with deeper architectures. The Table 3 below presents the F1 scores for each dataset, along with brief explanations of the modifications made in the deeper variants.The results show that the MLP embedding strategy consistently achieves the highest F1 scores for both datasets, indicating that the feed-forward architecture performs robustly compared to the other methods. Additionally, while the deeper versions of ResNorm and MLP exhibit comparable performance to their standard counterparts, the Transformer-based approaches record noticeably lower F1 scores, particularly on the Türkiye-S dataset.
@@ 3 Experiments
In this section, we evaluate the performance of the proposed PDA framework by conducting four experiments: (i) impact of integrating metadata with satellite imagery on the overall model performance, (ii) comparing the performance of the proposed framework with two previously published state-of-the-art machine learning frameworks, (iii) assessing the influence of individual metadata features on damage state prediction, and (iv) evaluating the effect of metadata in improving the trained model’s generalizability to different regions.
@@ 3.1 Impact of Adding Metadata on Performance: QuakeMetaFormer
Satellite imagery provides essential visual information for damage assessment, while earthquake-related metadata, such as PGA, SAR, VS30, etc., offers unique insights into event intensity and structural changes. However, relying on either data source individually often lacks the detail necessary for distinguishing between intermediate damage classes. The objective of this experiment is to evaluate how incorporating metadata with satellite imagery impacts the performance of the QuakeMetaFormer model. To achieve this, we conducted three experiments: (i) QIF-Türkiye-L, where the QuakeImageFormer (QIF) was trained using only satellite imagery to establish baseline performance, (ii) QMF-Türkiye-L, where the QuakeMetaFormer (QMF) model was trained using both satellite imagery and all the metadata, including aftershock seismic intensity indicators; and (iii) QMF-Türkiye-L1, where the QuakeMetaFormer model was trained using both satellite imagery and metadata, but with the seismic intensity indicators for aftershocks removed, retaining only data from the main event of the largest magnitude. Similarly, two additional experiments, QIF-Türkiye-S and QMF-Türkiye-S, were conducted on the Türkiye-S dataset. Comparing the results from these experiments allows us to understand the impact of adding metadata to satellite imagery in multiclass damage classification. Additionally, to evaluate the impact of metadata on the model’s ability to interpret imagery, we conducted four additional experiments. In these experiments, metadata was incorporated during training but masked during testing. Three experiments, QMF-Türkiye-L-M, QMF-Türkiye-L1-M and QMF-Türkiye-S-M, simulate scenarios where all metadata are masked during testing, while QMF-Türkiye-L1-M-SAR represents the case where only SAR-derived parameters (SAR-VV, SAR-VH, and DPM) are masked. This setup is particularly relevant for rapid post-earthquake damage assessment, where high-latency SAR data may be delayed. For example, Vitale and Milillo (2024) notes that for high-resolution SAR acquisitions in spotlight mode, coverage latency can extend up to 11 days for areas larger than 9,247 km², making it impractical for timely response over large disaster zones such as the  350,000 km² region affected by the 2023 Turkey–Syria earthquake. These experiments allow us to evaluate the practical applicability of QuakeMetaFormer when certain metadata are unavailable at inference time.
Different models trained are summarized in Table 4. In each of our experiments, we utilize 85% of the data for training and reserve 15% for the testing set to evaluate model performance. Each experiment is discussed in detail in the following sections.
@@
QMF-Türkiye-L QuakeMetaFormer (QMF) Türkiye-L Yes No
QIF-Türkiye-L QuakeImageFormer (QIF) Türkiye-L No N/A
QMF-Türkiye-L-M QuakeMetaFormer (QMF) Türkiye-L Yes Yes
QIF-Türkiye-L-KAH QuakeImageFormer (QIF) Türkiye-L No N/A
QMF-Türkiye-L1 QuakeMetaFormer (QMF) Türkiye-L Yes No
QMF-Türkiye-L1-M QuakeMetaFormer (QMF) Türkiye-L Yes Yes
QMF-Türkiye-L1-M-SAR QuakeMetaFormer (QMF) Türkiye-L Yes Yes
QMF-Türkiye-S QuakeMetaFormer (QMF) Türkiye-S Yes No
QMF-Türkiye-S-M QuakeMetaFormer (QMF) Türkiye-S Yes Yes
QIF-Türkiye-S QuakeImageFormer (QIF) Türkiye-S No N/A
    Tested only on Kahramanmaraş Region
@@ 3.2 Comparison of QuakeMetaFormer and Machine Learning Frameworks
To evaluate the performance of our proposed framework against other state-of-the-art PDA frameworks for earthquake damage assessment, we compared different architectures trained on the same dataset. For this experiment, we trained the QuakeMetaFormer model (QMF-Türkiye-S) on the Türkiye-S dataset (Imagery and metadata) and compared its performance with the best-performing machine learning model (random forest) reported by Yu et al. (2024), which was trained on the same metadata from the Türkiye-S dataset. On scanning the literature, we found another paper, Rao et al. (2023) that conducted a multiclass classification study for post-earthquake damage evaluation across regions impacted by 2015 Gorkha, 2017 Puebla, 2020 Puerto Rico, and 2020 Zagreb earthquakes. However, the datasets from Rao et al. (2023) were not accessible, so we report the average performance of our model across both our datasets and the average performance across all datasets in Rao et al. The experiments are summarized in Table 5.
@@
QMF-Türkiye-S QuakeMetaFormer (QMF) Our Study
RF-Yu Random Forest (RF) Yu et al. (2024)
@@ 3.3 Feature Importance Across Metadata
Various metadata provide different insights related to earthquake damage assessment. To assess the contribution of individual metadata features, we conducted a detailed analysis using SHapley Additive exPlanations (SHAP) Lundberg and Lee (2017). SHAP provides a robust method for understanding the output of machine learning models by perturbing each feature and measuring the resulting changes in prediction probabilities. These changes are then used to assign an importance score based on the feature’s contribution to the predictions and reveal how specific features increase the model’s confidence (probabilities) in predicting certain damage classes. For this experiment, we applied SHAP to the QMF-Türkiye-L1 model to evaluate the influence of various metadata inputs on damage state prediction. This analysis helps in understanding the impact of different earthquake-related indices on model performance. We also performed a class-wise analysis of metadata features to better understand the impact on different classes.